Derive the expression for the energy stored by a capacitor of charge Q potential difference between the plate of V and capacitance.
1
Expert's answer
2020-11-23T10:31:05-0500

Let there be a capacitor with charge "Q" , potential difference across it "V" and it's capacitance "C."


We have "Q=CV"


Also work done "W=VQ"


If the source delivers a small amount of charge "dQ" at a constant potential "V" , the work done "dW=dQ\times V =\frac{Q}{C}dQ"


Now the total work to deliver a charge "Q" is given by


"W=\displaystyle\int_{0}^{Q} \frac{Q}{C}dQ=\frac{Q^2}{2C}"


Work done =energy stored in the capacitor.


Substituting "Q=CV" in above equation we have,


Energy stored in capacitor "= E=\frac{Q^2}{2C}=\frac{1}{2} CV^2"